A COMPLETE STRUCTURE OF TRANS-3,4-DIFLUOROCYCLOBUTENE FROM MICROWAVE SPEC-TROSCOPY

\underline{NORMAN C. CRAIG}, LEWIS V. MCCARTY, PETER T. LINGENFELTER, A. SONAN OS-MANI, and OSMAN RATHORE, Department of Chemistry, Oberlin College, Oberlin, OH 44074; MICHAEL J. TUBERGEN, Department of Chemistry, Kent State University, Kent, OH 44242.; ROBERT L. KUCZKOWSKI, Department of Chemistry, University of Michigan, Ann Arbor, MI 48109.

Microwave spectra have been observed in the 6-17-GHz region for trans-3,4-difluorocyclobutene, two ^{13}C_{1} modifications, two d_{1} modifications, and the perdeutero species. Watson-type Hamiltonians have been fit with a full set of quartic centrifugal distortion constants for all species except for the d_{4} species. Rotational constants for the normal species in MHz are A = 5664.02410(29), B = 3366.07000(19), and C = 2345.58058(16). Complete structures fit with an r_{s}/r_{0} treatment and an r_{0} treatment are in satisfactory agreement. Preferred r_{s}/r_{0} results with Costain uncertainties in parentheses are 1.349(4) Å for the C=C bond, 1.503(9) Å for the contiguous C-C bonds, 1.534(4) Å for the unique C-C bond, 1.398(6) Å for the C-F bond, 1.081(3) Å for the =C-H bond, and 1.102(5) Å for the -C-H bond. Consistent with C_{2} symmetry, the ring is slightly puckered. Compared with cyclobutene, the C=C bond is slightly longer and the C-C bonds are shorter, adjustments which are familiar consequences of fluorine substitution on small rings.